Directions: Solve the equation by completing the square. Express any complex numbers using i notation.
x^2+12x=-24

Answer by Nate(3500)

x^2 + 12x = -24
(x + 6)^2 = -24 + 36
(x + 6)^2 = 12
x + 6 = +- sqrt(12)
x = -6 +- 2sqrt(3)